# Inference for Regression with

### v2

@@ 2.2 Two-Step Strategy
However, Proposition 1 also shows that two-step inference is invalid when $\kappa>0$. In this case, $\hat{\gamma}_{1}$ is consistent and its asymptotic variance is the same as if $Y_{i}$ were regressed on the true $\theta_{i}$, but the center of the asymptotic distribution is shifted due to the effect of measurement error. Confidence intervals based on standard OLS inference will therefore have approximately correct width but incorrect centering, meaning that their coverage rates will be below nominal coverage. [^9] It follows from the general treatment in Section 3 that Eicker–Huber–White standard errors based on the estimated $\hat{\theta}_{i}$ are consistent.
@@ 2.3 Upstream Inference
on $Z_{i}$. Let $\hat{\phi}_{1}$ denote the corresponding OLS estimator. To understand the forces at play, we study the behavior of $\hat{\phi}_{1}$ in a sequence of populations where the distribution of $(X_{i},Y_{i},Z_{i},\theta_{i})$ is fixed but the distribution of $C_{i}$ varies with $n$ so that (3) holds. Like before, to facilitate derivations we assume $(X_{i},C_{i})$ and $Z_{i}$ are independent conditional on $\theta_{i}$, and $C_{i}$ and $(\theta_{i},Z_{i})$ are independent.
@@ 3 Full Analysis of the Two-Step Strategy
One limiting feature of the stylized model in the previous section is that the observed data are not high-dimensional. In this section, we allow each unstructured observation to lie in a high-dimensional space, which requires some dimensionality reduction prior to regression analysis. We first describe the statistical framework linking unstructured data and the downstream regression model. We then analyze the two-step strategy and show why it leads to biased inference in empirically plausible settings. In the next section, we discuss how the one-step strategy solves these inference problems.
@@ 3.1 Statistical Framework
We begin by specifying a statistical model that, broadly speaking, has two parts. The first part computes low-dimensional numerical representations of the unstructured data. The second part introduces these numerical representations as covariates, potentially along with other quantitative data, into a linear regression model. There are a wide array of methods for dimensionality reduction used in the literature. We focus on factor modeling, which, in the context of high-dimensional discrete data, is also called topic modeling. We make this choice for two reasons. First, topic models have a well-defined statistical structure which facilitates theoretical analysis. Second, there is a large empirical literature which uses topic models as part of the two-step strategy, most commonly using textual data. Examples include Hansen et al., (2018), Mueller and Rauh, (2018), Larsen and Thorsrud, (2019), Thorsrud, (2020), Bybee et al., (2020), Adams et al., (2021), and Ash et al., (2022). However, the use of topic models is not limited to textual data. For example, Bandiera et al., (2020), Draca and Schwarz, (2021), and Munro and Ng, (2022) use topic models to analyze survey data. Meanwhile, Nimczik, (2017) and Olivella et al., (2021) use topic models for network-structured data.
@@ 3.1.1 Model
Many types of unstructured data are high-dimensional and discrete. We therefore let each unstructured observation $i$ be described by $\mathbf{x}_{i}$, a $V$-dimensional vector of count variables. The value $x_{i,v}$ is the number of times a feature $v$ appears in observation $i$. We allow $V$ to be high dimensional. For example, in the bag-of-words model $V$ is the number of unique terms in a textual corpus, typically in the thousands, and $x_{i,v}$ is the count of term $v$ in document $i$.
@@
The model we consider for the unstructured data is widely used in practice and tractable enough that we can develop clean characterizations for the two-step strategy. As $\mathbf{x}_{i}$ is a vector of counts, it is without loss of generality to model it as Multinomial. We impose some structure on the count probabilities for interpretability. The model is based on Probabilistic Latent Semantic Analysis (Hofmann, 1999, PLSA), a widely used factor model for discrete data, and its close cousin Latent Dirichlet Allocation (Blei et al., 2003, LDA). Formally,
@@
where $C_{i}=\sum_{v=1}^{V}x_{i,v}$ is the count of all features in observation $i$—a measure of the amount of unstructured data for observation $i$—and the count probabilities $\mathbf{p}_{i}$ have the factor structure $\mathbf{p}_{i}=\mathbf{B}^{T}\bm{\theta}_{i}$. This model nests as a special case a pure multinomial model where $\mathbf{B}=\mathbf{I}$ and $\bm{\theta}_{i}=\mathbf{p}_{i}$.
There are $K$ separate distributions over the $V$ features denoted $\bm{\beta}_{1},\ldots,\bm{\beta}_{K}$ where each $\bm{\beta}_{k}$ lies in the $(V-1)$-dimensional simplex. The distributions are called topics in text applications. More generally, they represent common factors from which individual observations are built. We collect the factors into a $K\times V$ row-stochastic matrix $\mathbf{B}$ where $\mathbf{B}^{T}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{K}]$. Each observation $i$ is characterized by the latent vector $\bm{\theta}_{i}$ which lies in the $(K-1)$-dimensional simplex. Its elements $\theta_{i,k}$ represent the weight attached to $\bm{\beta}_{k}$ in generating $\mathbf{x}_{i}$. Hence, the count probabilities for observation $i$ are $\mathbf{p}_{i}=\sum_{k=1}^{K}\bm{\beta}_{k}\theta_{i,k}$. Note that $\mathbf{B}$ is a matrix of common parameters whereas $\bm{\theta}_{i}$ is an observation-specific latent random vector. The quantity $C_{i}$ determines the degree of precision with which we can infer $\bm{\theta}_{i}$ from $\mathbf{x}_{i}$. The interplay between $C_{i}$ and the number of observations $n$ plays an important role in our theory.
@@
Suppose each unstructured observation is a monetary policy speech. One distribution $\bm{\beta}_{k}$ might put high weight on words like ‘inflation’, ‘prices’, and ‘cpi’, so $\bm{\beta}_{k}$ would have an interpretation as price rises. The corresponding $\theta_{i,k}$ represents how much speech $i$ discuss price rises. One research question might ask how attention paid to price rises, along with other economic conditions captured by other topics, affects policy actions. This could be captured by the $\bm{\gamma}$ coefficients in (4) where $Y_{i}$ is the policy action of speaker $i$, $\bm{S}$ selects the price rises topic weight from $\bm{\theta}_{i}$ and discards irrelevant topics (e.g., words used in generic conversation), and $\mathbf{q}_{i}$ measures quantitative information like market forecasts for growth and inflation at the time the speech was made.
@@ 3.1.2 Data and Maintained Assumptions
The data are a random sample $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ satisfying (4) and (5). To simplify derivations, we further assume that $C_{i}$ is independent of $(\bm{\theta}_{i},\mathbf{q}_{i},Y_{i})$, and that $(\mathbf{q}_{i},Y_{i})$ and $\mathbf{x}_{i}$ are independent conditional on $(C_{i},\bm{\theta}_{i})$. We also assume that $\mathbf{B}$ is identified. That is, there is a unique decomposition $\mathbf{P}=\mathbf{B}^{T}\bm{\Theta}$ with $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{n}]$ collecting the vectors of count probabilities across observations and $\bm{\Theta}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{n}]$ collecting the topic weights across observations. w Identification is commonly achieved in text applications by assuming the existence of anchor words that are known to appear in some topics but not others. We assume identifiability because our objective is to analyze the consequences of the two-step inference approach in a transparent way. Adding partial identification will significantly complicate the analysis but may be an interesting extension in future research.
@@ 3.2 Theory for the Two-Step Strategy
Evidently there is a measurement error problem: the estimates $\hat{\bm{\theta}}_{i}$ are noisy proxies for the true $\bm{\theta}_{i}$ appearing in the regression model (4). But Step (ii) overlooks this problem and treats the first-stage estimates $\hat{\bm{\theta}}_{i}$ as regular numeric data. This raises the possibility that OLS estimates of $\bm{\psi}$ may be biased due to measurement error introduced in Step (i). Moreover, conventional standard errors are typically reported. But these do not account for any additional variation introduced by using $\hat{\bm{\theta}}_{i}$ instead of $\bm{\theta}_{i}$.
@@
$\bm{\psi}=\left[\begin{array}[]{c}\bm{\gamma}\\ \bm{\alpha}\end{array}\right],\quad\quad\bm{\xi}_{i}=\left[\begin{array}[]{c}\bm{S}\bm{\theta}_{i}\\ \mathbf{q}_{i}\end{array}\right],\quad\quad\hat{\bm{\xi}}_{i}=\left[\begin{array}[]{c}\bm{S}\hat{\bm{\theta}}_{i}\\ \mathbf{q}_{i}\end{array}\right].$
The OLS estimator of $\bm{\psi}$ in the two-step strategy is given by
$\hat{\bm{\psi}}=\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\right).$ (6)
@@ 3.2.1 Fixed Population
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{2}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
@@
$\hat{\bm{\psi}}\to_{p}\bigg{(}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]+\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left(\mathbf{Q}_{\mathbf{B}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\bigg{)}^{-1}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\,\bm{\psi}\,.$
@@
$\bm{b}=\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\left[\begin{array}[]{c}\bm{S}\left(\mathbf{Q}_{\mathbf{B}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$
@@ 3.2.2 Sequence of Populations
We now consider a sequence of populations where the amount of unstructured data per observation becomes larger as the sample size $n$ increases. This asymptotic framework is designed to shed light on how $\hat{\bm{\psi}}$ behaves in empirically realistic settings where there are a relatively large number of observations and there is a large amount of unstructured data per observation. In this scenario, the measurement errors for each observation are small but their cumulative effect may not be ignorable relative to sampling error.
Formally, we consider a sequence of populations indexed by sample size $n$. In each population, we keep the distribution of $(Y_{i},\bm{\theta}_{i},\mathbf{q}_{i})$ conditional on $C_{i}$ fixed and as described in Section 3.1. We also maintain the assumption that $\mathbf{x}_{i}$ is given by the topic model (5). However, we let the marginal distribution of $C_{i}$ change with the sample size $n$ to allow the amount of unstructured data per observation to become large as the sample size $n$ increases. Specifically, we consider a framework in which
@@
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{4}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
@@
Assumption 2(i) is the same as Assumption 1(i). Assumption 2(ii)-(iii) strengthens Assumption 1(ii)-(iii) to require convergence a faster-than-root-$n$ rate. We believe Assumption 2(ii) is broadly satisfied in view of known convergence rates for estimators of $\mathbf{B}$ [^11] Bing et al., (2020), Wu et al., (2023), and Ke and Wang, (2022) derive finite-sample guarantees for different estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$. Each of their results implies the corresponding estimator $\hat{\mathbf{B}}$ converges at the optimal rate $(nC)^{-1/2}$ (up to log terms) where, for simplicity, the $C_{i}$ are all of the same order $C$. Hence, all estimators $\hat{\mathbf{B}}$ converge faster than $n^{-1/2}$ when $C$ grows with $n$, as we have here by (7). Assumption 2(iii) is made to simplify derivations but is not vacuous: given any estimator $\hat{\mathbf{B}}$, one could set $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$, in which case $\sqrt{n}\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. As before, Assumption 2(i)-(iii) holds trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$. Assumptions 2(iv) and 2(v) are standard. Assumption 2(vi) is made to simplify technical derivations and can be relaxed. It implies that $C_{i}$ is supported on $[c(\log n)^{1+\epsilon},\infty)$ for some $c,\epsilon>0$. This is weaker than the conventional assumption that all $C_{i}$ grow at the same rate $C$ (Bing et al., 2020, Wu et al., 2023, Ke and Wang, 2022) which, in view of (7), would imply that $C_{i}$ is supported on $[cn^{1/2},\infty)$. This condition is only used to establish consistency of standard errors.
@@
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\\ \to_{d}N\left(\kappa\bm{b},\,\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right),$ (8)
@@
$\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}.$ (9)
@@
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right).$

### v3

@@ 2.2 Two-Step Strategy
However, Proposition 1 also shows that two-step inference is invalid when $\kappa>0$. In this case, $\hat{\gamma}_{1}$ is consistent and its asymptotic variance is the same as if $Y_{i}$ were regressed on the true $\theta_{i}$, but the center of the asymptotic distribution is shifted due to the effect of measurement error. Confidence intervals based on standard OLS inference will therefore have approximately correct width but incorrect centering, meaning that their coverage rates will be below nominal coverage. [^7] It follows from the general treatment in Section 3 that Eicker–Huber–White standard errors computed from $\hat{\theta}_{i}$ instead of $\theta_{i}$ are consistent.
@@ 2.3 Upstream Inference
on $Z_{i}$. Let $\hat{\phi}_{1}$ denote the corresponding OLS estimator. To understand the forces at play, we study the behavior of $\hat{\phi}_{1}$ in a sequence of populations where the distribution of $(X_{i},Y_{i},Z_{i},\theta_{i})$ conditional on $C_{i}$ is fixed but the distribution of $C_{i}$ varies with $n$ so that (3) holds. Like before, to facilitate derivations we assume $(X_{i},C_{i})$ and $Z_{i}$ are independent conditional on $\theta_{i}$, and $C_{i}$ and $(\theta_{i},Z_{i})$ are independent.
@@ 3 Full Analysis of the Two-Step Strategy
One limiting feature of the stylized model in the previous section is that the observed data are not high-dimensional. In this section, we allow each unstructured observation to lie in a high-dimensional space, which requires some dimensionality reduction prior to regression analysis. We first describe the statistical framework linking unstructured data and the downstream regression model. We then analyze the two-step strategy and show why it leads to biased inference in empirically plausible settings.
@@ 3.1 Statistical Framework
We begin by specifying a statistical model that, broadly speaking, has two parts. The first computes low-dimensional numerical representations of the unstructured data. The second introduces these numerical representations as covariates, potentially along with other quantitative data, into a linear regression model. There is a wide array of methods for dimensionality reduction used in the literature. We focus on factor modeling, which, in the context of high-dimensional discrete data, is also called topic modeling. We make this choice for two reasons. First, topic models have a well-defined statistical structure which facilitates theoretical analysis. Second, there is a large empirical literature which uses topic models as part of the two-step strategy, most commonly using textual data. Examples include Hansen et al., (2018), Mueller and Rauh, (2018), Larsen and Thorsrud, (2019), Thorsrud, (2020), Bybee et al., (2020), Adams et al., (2021), and Ash et al., (2022). However, the use of topic models is not limited to textual data. For example, Bandiera et al., (2020), Draca and Schwarz, (2021), and Munro and Ng, (2022) use topic models to analyze survey data. Meanwhile, Nimczik, (2017) and Olivella et al., (2021) use topic models for network-structured data.
@@ 3.1.1 Model
Many types of unstructured data are high-dimensional and discrete. We therefore let each unstructured observation $i$ be described by $\mathbf{x}_{i}=(x_{i,v})_{v=1}^{V}$, a $V$-dimensional vector of count variables. The value $x_{i,v}$ is the number of times a feature $v$ appears in observation $i$. For example, in the bag-of-words model $V$ is the number of unique terms in a textual corpus, typically in the thousands, and $x_{i,v}$ is the count of term $v$ in document $i$.
@@
The model we consider for the unstructured data is widely used in practice and tractable enough that we can develop clean characterizations for the two-step strategy. As $\mathbf{x}_{i}$ is a vector of counts, it is without loss of generality to model it as Multinomial. We impose some structure on the count probabilities for interpretability. The model is based on Probabilistic Latent Semantic Analysis (Hofmann, 1999), a widely used factor model for discrete data, and its close cousin Latent Dirichlet Allocation (Blei et al., 2003, LDA).
There are $K$ separate distributions over the $V$ features denoted $\bm{\beta}_{1},\ldots,\bm{\beta}_{K}$ where each $\bm{\beta}_{k}$ lies in the $(V-1)$-dimensional simplex. The distributions are called topics in text applications. More generally, they represent common factors from which individual observations are built. We collect the factors into a $K\times V$ row-stochastic matrix $\mathbf{B}$ where $\mathbf{B}^{T}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{K}]$. Each observation $i$ is characterized by the latent vector $\bm{\theta}_{i}$ which lies in the $(K-1)$-dimensional simplex. Its elements $\theta_{i,k}$ represent the weight attached to $\bm{\beta}_{k}$ in generating $\mathbf{x}_{i}$. Hence, the count probabilities for observation $i$ are $\mathbf{p}_{i}=\sum_{k=1}^{K}\bm{\beta}_{k}\theta_{i,k}$. Note that $\mathbf{B}$ is a matrix of common parameters whereas $\bm{\theta}_{i}$ is an observation-specific latent random vector. Putting these elements together, the distribution of $\mathbf{x}_{i}$ can be written as
@@
where $C_{i}=\sum_{v=1}^{V}x_{i,v}$ is the count of all features in observation $i$—a measure of the amount of unstructured data for observation $i$—and the count probabilities $\mathbf{p}_{i}$ have the factor structure $\mathbf{p}_{i}=\mathbf{B}^{T}\bm{\theta}_{i}$. This model nests as a special case a pure multinomial model where $K=V$, $\mathbf{B}=\mathbf{I}$, and $\bm{\theta}_{i}=\mathbf{p}_{i}$. The quantity $C_{i}$ determines the degree of precision with which we can infer $\bm{\theta}_{i}$ from $\mathbf{x}_{i}$. The interplay between $C_{i}$ and the number of observations $n$ plays an important role in our theory.
@@
Suppose each unstructured observation is a monetary policy speech. One distribution $\bm{\beta}_{k}$ might put high weight on words like ‘inflation’, ‘prices’, and ‘cpi’, so $\bm{\beta}_{k}$ would have an interpretation as price rises. The corresponding $\theta_{i,k}$ represents how much speech $i$ discusses price rises. One research question might ask how attention paid to price rises, along with other economic conditions captured by other topics and numeric data, affects policy actions. This could be captured by the $\bm{\gamma}$ coefficients in (4) where $Y_{i}$ is the policy action of speaker $i$, $\bm{S}$ selects the price rises topic weight from $\bm{\theta}_{i}$ and discards irrelevant topics (e.g., words used in generic conversation), and $\mathbf{q}_{i}$ measures quantitative information like market forecasts for growth and inflation at the time the speech was made.
@@ 3.1.2 Data and Maintained Assumptions
The data are a random sample $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ satisfying (4) and (5). To simplify derivations, we further assume that $C_{i}$ is independent of $(\bm{\theta}_{i},\mathbf{q}_{i},Y_{i})$, and that $(\mathbf{q}_{i},Y_{i})$ and $\mathbf{x}_{i}$ are independent conditional on $(C_{i},\bm{\theta}_{i})$. We also assume that $\mathbf{B}$ is identified. That is, there is a unique decomposition $\mathbf{P}=\mathbf{B}^{T}\bm{\Theta}$ with $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{n}]$ collecting the vectors of count probabilities across observations and $\bm{\Theta}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{n}]$ collecting the topic weights across observations. Identification is commonly achieved in text applications by assuming the existence of anchor words that are known to appear in some topics but not others. We assume identifiability because our objective is to analyze the consequences of the two-step inference approach in a transparent way. Adding partial identification will significantly complicate the analysis but may be an interesting extension in future research.
@@ 3.2 Theory for the Two-Step Strategy
Evidently there is a measurement error problem: the estimates $\hat{\bm{\theta}}_{i}$ are noisy proxies for the true $\bm{\theta}_{i}$ appearing in the regression model (4). But Step (ii) overlooks this problem and treats the first-stage estimates $\hat{\bm{\theta}}_{i}$ as regular numeric data. This raises the possibility that OLS estimates may be biased due to measurement error introduced in Step (i). Moreover, conventional standard errors are typically reported. But these do not account for any additional variation introduced by using $\hat{\bm{\theta}}_{i}$ instead of $\bm{\theta}_{i}$.
@@
$\bm{\xi}_{i}=\left[\begin{array}[]{c}\bm{S}\bm{\theta}_{i}\\ \mathbf{q}_{i}\end{array}\right],\quad\quad\hat{\bm{\xi}}_{i}=\left[\begin{array}[]{c}\bm{S}\hat{\bm{\theta}}_{i}\\ \mathbf{q}_{i}\end{array}\right].$
The OLS estimator of $\bm{\psi}=\left[\bm{\gamma},\bm{\alpha}\right]^{T}$ in the two-step strategy is given by
$\hat{\bm{\psi}}=\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\right).$ (6)
@@ 3.2.1 Fixed Population
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{2}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
@@
$\hat{\bm{\psi}}\to_{p}\bigg{(}\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]+\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left(\mathbf{Q}_{\mathbf{B}}^{\phantom{T}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\bigg{)}^{-1}\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\,\bm{\psi}\,.$
@@
$\bm{b}=\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\left[\begin{array}[]{c}\bm{S}\left(\mathbf{Q}_{\mathbf{B}}^{\phantom{T}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$
@@ 3.2.2 Sequence of Populations
We now consider a sequence of populations where the amount of unstructured data per observation becomes larger as the sample size $n$ increases. This asymptotic framework is designed to shed light on how $\hat{\bm{\psi}}$ behaves in empirically realistic settings where there are a relatively large number of observations and there is a large amount of unstructured data per observation. In this scenario, the measurement errors across observations are small but their cumulative effect may not be ignorable relative to sampling error.
Formally, we consider a sequence of populations indexed by sample size $n$. In each population, we keep the distribution of $(\mathbf{x}_{i},Y_{i},\bm{\theta}_{i},\mathbf{q}_{i})$ conditional on $C_{i}$ fixed and as described in Section 3.1. However, we let the marginal distribution of $C_{i}$ change with the sample size $n$ to allow the amount of unstructured data per observation to become large as the sample size $n$ increases. Specifically, we consider a framework in which
@@
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{4}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
@@
Assumption 2(i) is the same as Assumption 1(i). Assumption 2(ii)-(iii) strengthens Assumption 1(ii)-(iii) to require convergence at a faster-than-root-$n$ rate. We believe Assumption 2(ii) is broadly satisfied in view of known convergence rates for estimators of $\mathbf{B}$ [^9] Bing et al., (2020), Wu et al., (2023), and Ke and Wang, (2022) derive finite-sample guarantees for different estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$. Each of their results implies the corresponding estimator $\hat{\mathbf{B}}$ converges at the optimal rate $(nC)^{-1/2}$ (up to log terms) where, for simplicity, the $C_{i}$ are all of the same order $C$. Hence, all estimators $\hat{\mathbf{B}}$ converge faster than $n^{-1/2}$ when $C$ grows with $n$, as we have here by (7). Assumption 2(iii) is made to simplify derivations but is not vacuous: given any estimator $\hat{\mathbf{B}}$, one could set $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$, in which case $\sqrt{n}\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. As before, Assumption 2(i)-(iii) holds trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$. Assumptions 2(iv) and 2(v) are standard. Assumption 2(vi) is made to simplify technical derivations and can be relaxed. It implies that $C_{i}$ is supported on $[c(\log n)^{1+\epsilon},\infty)$ for some $c,\epsilon>0$. This is weaker than the conventional assumption that all $C_{i}$ grow at the same rate $C$ (Bing et al., 2020, Wu et al., 2023, Ke and Wang, 2022) which, in view of (7), would imply that $C_{i}$ is supported on $[cn^{1/2},\infty)$ for some $c>0$. This condition is only used to establish consistency of standard errors.
@@
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\\ \to_{d}N\left(\kappa\bm{b},\,\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right),$ (8)
@@
$\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}.$ (9)
@@
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right).$